# Predictability Enables Parallelization of Nonlinear State Space Models

### v2

@@ Appendix C Merit Function PL Constant is Controlled by Largest Lyapunov Exponent of Model
Assume that the LLE regularity condition (10) holds. Then if $\lambda\neq 0$ the PL constant $\mu$ of the merit function in (8) satisfies
@@
By L’Hôpital’s rule, if $\lambda=0$, then the bounds are instead
$\frac{1}{aT}\ \leq\ \sqrt{\mu}\ \leq\ \frac{1}{b}.$
@@
the result follows by applying eq.˜16 and eq.˜19 at all $\mathbf{s}^{(i)}$ along the optimization trajectory. ∎
@@
The LLE regularity conditions in eq.˜10 highlight the more natural “average case” behavior experienced along actual trajectories $\mathbf{s}\in\mathbb{R}^{TD}$. This “average case” behavior is highlighted, for example, by our experiments with the two-well system (cf. Section 5 and Appendix J.4), where even though a global upper bound on $\|J_{t}(s_{t})\|_{2}$ over all of state space would be greater than $1$ (i.e., there are unstable regions of state space), we observe fast convergence of DEER because the system as a whole has negative LLE (its trajectories are stable on average).
@@
However, because our bounds in eq.˜22 and eq.˜23 on $\sigma_{\max}(\mathbf{J})$ do not scale with the sequence length $T$, it follows that the scaling with $T$ of an upper bound on $\kappa(\mathbf{J})$—the conditioning of the optimization problem—is controlled solely by the bounds on $\sigma_{\min}(\mathbf{J})$ that we provided in Theorem 2. The importance of studying how the conditioning scales with $T$ stems from the fact that we would like to understand if there are regimes—particularly involving large sequence lengths and parallel computers—where parallel evaluation can be faster than sequential evaluation.
@@ Appendix G Alternative Descent Techniques & Worst/Average Complexity
In practice, we observe that DEER converges much faster than the worst-case analysis (35) would suggest. In particular, we observe that DEER converges in roughly $\log\frac{1}{\mu}$, steps, even for unpredictable systems. This behavior can be explained with a simple “two-phase" model, wherein the DEER iterates move towards the basin of quadratic convergence at a rate which is independent of the PL-constant $\mu$ (see Appendix J.1).
@@ Appendix H Proof of Size of Basin of Quadratic Convergence
which is the direct analogy of Boyd and Vandenberghe [28, 9.33]. To reiterate, here $L$ is the Lipschitz constant of $\mathbf{J}$, while $\mu:=\inf_{i\in\mathbb{N}}\sigma^{2}_{\mathrm{min}}\left(\mathbf{J}(\mathbf{s}^{(i)})\right)$.
@@ Appendix I Parameterizing Nonlinear SSMs to be contractive
Appendix J Experimental Details and Discussion
@@
J.1 Deriving the Empirical Scaling of DEER
@@
J.2 Details and Discussion for mean-field RNN experiment
@@
In the design of the weight matrix $W$, we follow Engelken et al. [29]. In particular, we draw each entry $W_{ij}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathcal{N}(0,\nicefrac{{g^{2}}}{{D}})$, where $g$ is a scalar parameter. We then set $W_{ii}=0$ for all $i$ (no self-coupling of the neurons). A key point of Engelken et al. [29] is that by scaling the single parameter $g$, the resulting RNN goes from predictable to chaotic behavior. While Engelken et al. [29] computes the full Lyapunov spectrum in the limit $D\to\infty$, for finite $D$ we can compute a very accurate numerical approximation to the LLE (cf. Appendix J.6). In Figure 4, we verify numerically that there is a monotonic relationship between $g$ and the LLE of the resulting system, and that the min-max range for 20 seeds is small.
Figure 4: Robust relationship in mean field RNN between variance parameter $g$ and LLE of the system. For 20 seeds, we observe a robust and monotonic relationship between the scalar parameter $g$ and the LLE of the resulting mean-field RNN. The plot above is made for $50$ different values of $g$ from $0.5$ to $2.0$ (linearly spaced).
@@
First, as discussed in Gonzalez et al. [3], DEER can experience numerical overflow when deployed on unstable systems. While we reset to the initialization (in this experiment we initialized $\mathbf{s}_{1:T}$ with iid draws from $\mathcal{U}[0,1]$), doing so slows convergence. Thus, many of our runs for $\lambda>0$ and large $T$ take the maximum number of DEER iteration we allow (we do not allow more than $T$ iterations, as this is the theoretical upper bound for number of DEER iterations before convergence, cf. Proposition 1 of [3]), which helps to explain the slight increase in red space for experiment (center plot of Figure 2) vs. theory (left plot of Figure 2). Note, however, that for $T=1000$ (the sequence length shown in the right plot), there is no numerical overflow for the DEER trajectories for any of the 20 random seeds or 50 values of $g$ tried.
@@
J.3 Additional experiment for the mean-field RNN: other optimizers and wallclock time
@@
We include a quasi-Newton algorithm proposed in Gonzalez et al. [3] called quasi-DEER. Quasi-DEER simply replaces the $J_{t}$ defined in eq.˜4 with $\mathrm{diag}(J_{t})$, and so is also parallelizable over the sequence length with a parallel scan. Furthermore, we also include gradient descent on the merit function, which is embarrassingly parallel over the sequence length. In the top panel of Figure˜6, we observe that the number of steps for gradient descent and quasi-DEER to converge also scales monotonically with the LLE, as we expect from ˜2. DEER (Gauss-Newton) converges in a small number of steps all the way up to the threshold between predictability and unpredictability ($\lambda=0$). Intuitively, the performance of the other optimizers degrades more quickly as unpredictability increases because quasi-Newton and gradient descent use less information about the curvature of the loss landscape.
@@
We note that sequential evaluation of eq.˜1 can also be thought of as block coordinate descent on the merit function $\mathcal{L}(\mathbf{s})$, where the block $s_{t}\in\mathbb{R}^{D}$ is optimized at optimization step $(t)$. The optimization of each block is a convex problem: simply minimize $\|s_{t}-f(s_{t-1}^{*})\|_{2}^{2}$, or equivalently set $s_{t}=f(s_{t-1}^{*})$. As sequential evaluation will always take $T$ steps to converge, we do not include it in the top panel of Figure˜6.
@@
Wallclock time
In the bottom panel of Figure˜6, we also report the wallclock times for these algorithms to run (our experiments are run on an H100 with 80 GB onboard memory). We observe that the run time of sequential evaluation (green) is effectively constant with respect to $\lambda$. We observe that in the predictable setting, DEER is an order of magnitude faster than sequential evaluation, while in the unpredictable regime, DEER is 1-2 orders of magnitude slower than sequential evaluation. This importance of using parallel evaluation only in predictable settings is a core practical takeaway from our theoretical contributions.
@@
We run the experiment in Figure˜6 on a smaller scale than the experiment in Figure˜2 (Right). In Figure˜6, we consider 5 random seeds for 16 values of $g$ equispaced between $0.5$ and $2.0$. Each wallclock time reported is the average of 5 runs for the same seed. We use a batch size of 1. While DEER (Gauss-Newton) and quasi-DEER effectively do not have a step size (they use a step size of $1$ always). For each value of $g$, we ran gradient descent with the following set of step sizes $\alpha$: $0.01,0.1,0.25,0.5,0.6,0.7,0.8,0.9,\text{and }1.0$. For each value of $g$, we then pick the step size $\alpha$ that results in the fastest convergence of gradient descent. For the smallest value of $g=0.5$, we use $\alpha=0.6$; for $g=0.6$, we use $\alpha=0.5$; and for all other values of $g$, we use $\alpha=0.25$. Future work may investigate more adaptive ways to tune the step size $\alpha$, or to use a learning rate schedule.
@@
J.4 Additional details for the two-well potential
@@
where $\epsilon$ is the step size and $w_{t}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathcal{N}(0,I_{D})$.  [^3] Notice that this is a discretization (with time step $\epsilon$) of the Langevin Diffusion SDE $ds(t)=-\nabla\phi(s(t))dt+\sqrt{2}dw(t)$, where $w(t)$ is Brownian motion [82, 83, 84]. Accordingly, the Jacobians of the dynamics (those used in DEER) take the form
@@
J.5 Building Stable Observers for Chaotic Systems
@@
J.6 Numerical computation of the discrete-time LLE
@@
Appendix K Discrete and Continuous Time LLE
@@
$\text{continuous-time:}\quad\quad\dot{x}(t)$ $=\dfrac{\partial F}{\partial x}\left(x(t),t\right)x(t).$
@@
All of our experiments are the forward Euler discretizations of continuous-time systems. So, in this section, we we work out what happens to the LLE in such a setting. Let’s consider running our continuous-time setting for a length of time $T$.
@@
Note that naively plugging in this discrete-time estimator would therefore result in a different estimate of the LLE depending on the size of the time-step $\Delta t$. For this reason, in our experiments we report an estimate of the continuous-time LLE $\lambda_{c}$ given by $\nicefrac{{\lambda_{d}(\Delta t)}}{{\Delta t}}$.
Note that dividing by $\Delta t$, which is positive, does not change the sign of $\lambda$, i.e. whether or not the system is predictable or unpredictable. However, estimating $\lambda_{c}$ has the benefit of giving stability to the estimate even as we change $\Delta t$, and allowing for comparison to the known LLEs of continuous time systems.

### v3

@@ Appendix C Merit Function PL Constant is Controlled by Largest Lyapunov Exponent of Model
Assume that the LLE regularity condition from eq.˜10 holds. Then if $\lambda\neq 0$ the PL constant $\mu$ of the merit function in (8) satisfies
@@
If $\lambda=0$, then the bounds are instead
$\frac{1}{aT}\ \leq\ \sqrt{\mu}\ \leq\ \frac{1}{b}\sqrt{\frac{2D}{T+1}}.$
@@
the result for $\lambda\neq 0$ follows by applying eq.˜16 and eq.˜19 at all $\mathbf{s}^{(i)}$ along the optimization trajectory.
Finally, let us conclude by considering the case $\lambda=0$. In this setting, the lower bound on $\sqrt{\mu}$ follows from L’Hôpital’s rule. For the upper bound, we again must lower bound $\|\mathbf{J}^{-1}\|_{2}$. To do so, we leverage the relationship between spectral and Frobenius norms, namely that for an $n\times n$ matrix $A$,
$\frac{\|A\|_{F}}{\sqrt{n}}\leq\|A\|_{2}\leq\|A\|_{F}.$ (20)
We can find the squared Frobenius norm, i.e. $\|\mathbf{J}^{-1}\|_{F}^{2}$, which is the sum of the squares of all of the entries. The squared Frobenius norm factors over the block structure of the matrix, i.e. $\|\mathbf{J}^{-1}\|_{F}^{2}$ is the sum of the squared Frobenius norms of the blocks. We know that each block has spectral norm lower bounded by $b$, so each block also has Frobenius norm lower bounded by $b$. Therefore, summing up over all of the blocks, it follows that
$b^{2}\frac{T(T+1)}{2}\leq\|\mathbf{J}^{-1}\|_{F}^{2}$
and
$\|\mathbf{J}^{-1}\|_{F}\leq\sqrt{TD}\|\mathbf{J}^{-1}\|_{2}.$
Putting these equations together, it follows that
$b\sqrt{\frac{T(T+1)}{2}}\leq\sqrt{TD}\|\mathbf{J}^{-1}\|_{2}$
or
$b\sqrt{\frac{T+1}{2D}}\leq\|\mathbf{J}^{-1}\|_{2},$
and so the upper bound on $\sqrt{\mu}$ when $\lambda=0$ follows from taking reciprocals. ∎
@@
Connection to semiseparable matrices and Mamba2
Having depicted the structure of $\mathbf{J}^{-1}$, we note the connection between $\mathbf{J}^{-1}$ in this paper and the attention or sequence mixer matrix $M$ in Dao and Gu [14], which introduced the Mamba2 architecture (see equation 6 or Figure 2 of Dao and Gu [14] for the form of $M$, and compare with $\mathbf{J}^{-1}$ above).
Mamba2 is a deep learning sequence modeling architecture. Its sequence mixer in each layer has at its core a linear dynamical system. Dao and Gu [14] observe that while a linear dynamical system (LDS) can be evaluated recurrently (sequentially) or in parallel (for example, with a parallel scan), it can also be evaluated multiplying the inputs to the LDS by the matrix $M$. Since each DEER iteration is also a linear dynamical system, with the transition matrices given by $\{J_{t}\}_{t=2}^{T}$, it follows that $M$ in Dao and Gu [14] and $\mathbf{J}^{-1}$ in our paper are the same object, and so results about these objects from these two papers transfer.
In particular, we observe that, in the language from Dao and Gu [14], the $\mathbf{J}^{-1}$ we consider in this paper is $D$-semiseparable (see Definition 3.1 from Dao and Gu [14]). Thus, any efficient, hardware-aware algorithms and implementations developed for $D$-semiseparable matrices could also be applied to accelerating each iteration of DEER, though we note that Dao and Gu [14] focus on the 1-semiseparable setting, which they call a state space dual or SSD layer. In any case, using these connections to accelerate each iteration of DEER and related parallel Newton algorithms from a systems implementation perspective would be an interesting direction for future work.
@@
The LLE regularity conditions in eq.˜10 highlight the more natural “average case” behavior experienced along actual trajectories $\mathbf{s}\in\mathbb{R}^{TD}$. This “average case” behavior is highlighted, for example, by our experiments with the two-well system (cf. Section 5 and Appendix K.4), where even though a global upper bound on $\|J_{t}(s_{t})\|_{2}$ over all of state space would be greater than $1$ (i.e., there are unstable regions of state space), we observe fast convergence of DEER because the system as a whole has negative LLE (its trajectories are stable on average).
@@
However, because our bounds in eq.˜23 and eq.˜24 on $\sigma_{\max}(\mathbf{J})$ do not scale with the sequence length $T$, it follows that the scaling with $T$ of an upper bound on $\kappa(\mathbf{J})$—the conditioning of the optimization problem—is controlled solely by the bounds on $\sigma_{\min}(\mathbf{J})$ that we provided in Theorem 2. The importance of studying how the conditioning scales with $T$ stems from the fact that we would like to understand if there are regimes—particularly involving large sequence lengths and parallel computers—where parallel evaluation can be faster than sequential evaluation.
@@ Appendix G Alternative Descent Techniques & Worst/Average Complexity
In practice, we observe that DEER converges much faster than the worst-case analysis (36) would suggest. In particular, we observe that DEER converges in roughly $\log\frac{1}{\mu}$, steps, even for unpredictable systems. This behavior can be explained with a simple “two-phase" model, wherein the DEER iterates move towards the basin of quadratic convergence at a rate which is independent of the PL-constant $\mu$ (see Appendix K.1).
@@ Appendix H Proof of Size of Basin of Quadratic Convergence
which is the direct analogy of Boyd and Vandenberghe [5, 9.33]. To reiterate, here $L$ is the Lipschitz constant of $\mathbf{J}$, while $\mu:=\inf_{i\in\mathbb{N}}\sigma^{2}_{\mathrm{min}}\left(\mathbf{J}(\mathbf{s}^{(i)})\right)$.
@@ Appendix I Parameterizing nonlinear SSMs to be contractive
Appendix J Interpreting nonlinear SSMs as stacks of linear dynamical systems
Figure 4: Equivalence between a contractive nSSM and an $\mathcal{O}(\log T)$ stack of linear state-space models. Contractivity implies that the nonlinear dynamics can be decomposed into a hierarchy of $\mathcal{O}(\log T)$ layers of linear SSMs, or linear dynamics systems (LDS), each of which can be evaluated in $\mathcal{O}(\log T)$ time by a parallel scan.
As mentioned in our Discussion in Section˜6, an important implication of our results is that a contractive nSSM can be interpreted as a hierarchical composition of linear state-space layers (SSMs), or equivalently, linear dynamical system (LDS) layers. Each layer can be evaluated in $\mathcal{O}(\log T)$ time with a parallel scan, and the total number of layers required scales as $\mathcal{O}(\log T)$. This perspective shows that nonlinear temporal dependencies can be captured through a logarithmic-depth stacking of linear dynamics. Figure 4 provides a schematic illustration of this equivalence.
More explicitly, each iteration of DEER is given by the linear dynamical system
$s_{t+1}^{(i+1)}=f(s_{t}^{(i)})+J_{t+1}(s_{t}^{(i)})\left(s_{t}^{(i+1)}-s_{t}^{(i)}\right).$ (38)
Therefore, we can interpret each “iteration” $(i)$ of DEER as a sequence-mixing “layer” $(i)$, where the sequence-mixing layer is an input-dependent switching linear dynamical system, like in Mamba [36]. The inputs to “layer” $(i+1)$ is the state trajectory of the immediately preceding “iteration” or “layer” $(i)$. Because we prove that DEER converges linearly in ˜4, it follows that a contractive nSSM can be simulated in $\mathcal{O}(\log T)$ LDS layers of the form shown in eq.˜38, assuming the initial error grows polynomially in the sequence length.
Appendix K Experimental Details and Discussion
@@
K.1 Deriving the Empirical Scaling of DEER
@@
K.2 Details and Discussion for mean-field RNN experiment
@@
In the design of the weight matrix $W$, we follow Engelken et al. [19]. In particular, we draw each entry $W_{ij}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathcal{N}(0,\nicefrac{{g^{2}}}{{D}})$, where $g$ is a scalar parameter. We then set $W_{ii}=0$ for all $i$ (no self-coupling of the neurons). A key point of Engelken et al. [19] is that by scaling the single parameter $g$, the resulting RNN goes from predictable to chaotic behavior. While Engelken et al. [19] computes the full Lyapunov spectrum in the limit $D\to\infty$, for finite $D$ we can compute a very accurate numerical approximation to the LLE (cf. Appendix K.6). In Figure 5, we verify numerically that there is a monotonic relationship between $g$ and the LLE of the resulting system, and that the min-max range for 20 seeds is small.
Figure 5: Robust relationship in mean field RNN between variance parameter $g$ and LLE of the system. For 20 seeds, we observe a robust and non-decreasing relationship between the scalar parameter $g$ and the LLE of the resulting mean-field RNN. The plot above is made for $50$ different values of $g$ from $0.5$ to $2.0$ (linearly spaced). We estimate the LLE over a sequence length of $T=9999$.
@@
First, as discussed in Gonzalez et al. [34], DEER can experience numerical overflow when deployed on unstable systems. While we reset to the initialization (in this experiment we initialized $\mathbf{s}_{1:T}$ with iid draws from $\mathcal{U}[0,1]$), doing so slows convergence. Thus, many of our runs for $\lambda>0$ and large $T$ take the maximum number of DEER iterations we allow (we do not allow more than $T$ iterations, as this is the theoretical upper bound for number of DEER iterations before convergence, cf. Proposition 1 of [34]), which helps to explain the slight increase in red space for experiment (center plot of Figure 2) vs. theory (left plot of Figure 2). Note, however, that for $T=1000$ (the sequence length shown in the right plot), there is no numerical overflow for the DEER trajectories for any of the 20 random seeds or 50 values of $g$ tried.
@@
K.3 Additional experiment for the mean-field RNN: other optimizers and wallclock time
@@
We include a quasi-Newton algorithm proposed in Gonzalez et al. [34] called quasi-DEER. Quasi-DEER simply replaces the $J_{t}$ defined in eq.˜4 with $\mathrm{diag}(J_{t})$, and so is also parallelizable over the sequence length with a parallel scan. Furthermore, we also include gradient descent on the merit function, which is embarrassingly parallel over the sequence length. In the top panel of Figure˜7, we observe that the number of steps for gradient descent and quasi-DEER to converge also scales monotonically with the LLE, as we expect from ˜2. DEER (Gauss-Newton) converges in a small number of steps all the way up to the threshold between predictability and unpredictability ($\lambda=0$). Intuitively, the performance of the other optimizers degrades more quickly as unpredictability increases because quasi-Newton and gradient descent use less information about the curvature of the loss landscape.
@@
We note that sequential evaluation of eq.˜1 can also be thought of as block coordinate descent on the merit function $\mathcal{L}(\mathbf{s})$, where the block $s_{t}\in\mathbb{R}^{D}$ is optimized at optimization step $(t)$. The optimization of each block is a convex problem: simply minimize $\|s_{t}-f(s_{t-1}^{*})\|_{2}^{2}$, or equivalently set $s_{t}=f(s_{t-1}^{*})$. As sequential evaluation will always take $T$ steps to converge, we do not include it in the top panel of Figure˜7.
Wallclock time
In the bottom panel of Figure˜7, we also report the wallclock times for these algorithms to run (our experiments are run on an H100 with 80 GB onboard memory). We observe that the run time of sequential evaluation (green) is effectively constant with respect to $\lambda$. We observe that in the predictable setting, DEER is an order of magnitude faster than sequential evaluation, while in the unpredictable regime, DEER is 1-2 orders of magnitude slower than sequential evaluation. This importance of using parallel evaluation only in predictable settings is a core practical takeaway from our theoretical contributions.
@@
We run the experiment in Figure˜7 on a smaller scale than the experiment in Figure˜2 (Right). In Figure˜7, we consider 5 random seeds for 16 values of $g$ equispaced between $0.5$ and $2.0$. Each wallclock time reported is the average of 5 runs for the same seed. We use a batch size of 1. While DEER (Gauss-Newton) and quasi-DEER effectively do not have a step size (they use a step size of $1$ always). For each value of $g$, we ran gradient descent with the following set of step sizes $\alpha$: $0.01,0.1,0.25,0.5,0.6,0.7,0.8,0.9,\text{and }1.0$. For each value of $g$, we then pick the step size $\alpha$ that results in the fastest convergence of gradient descent. For the smallest value of $g=0.5$, we use $\alpha=0.6$; for $g=0.6$, we use $\alpha=0.5$; and for all other values of $g$, we use $\alpha=0.25$. Future work may investigate more adaptive ways to tune the step size $\alpha$, or to use a learning rate schedule.
@@
K.4 Additional details for the two-well potential
@@
where $\epsilon$ is the step size and $w_{t}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathcal{N}(0,I_{D})$. In our experiments, we use $\epsilon=0.01$.  [^3] Notice that this is a discretization (with time step $\epsilon$) of the Langevin Diffusion SDE $ds(t)=-\nabla\phi(s(t))dt+\sqrt{2}dw(t)$, where $w(t)$ is Brownian motion [39, 62, 42]. Accordingly, the Jacobians of the dynamics (those used in DEER) take the form
@@
Futhermore, in this experiment we show empirically that Langevin dynamics can have negative LLE (cf. Figure˜3). This results suggest that the Metropolis-adjusted Langevin algorithm (MALA), a workhorse of MCMC, may also be predictable in settings of interest, including multimodal distributions.
@@
K.5 Building Stable Observers for Chaotic Systems
@@
K.6 Numerical computation of the discrete-time LLE
@@
Appendix L Discrete and Continuous Time LLE
@@
$\text{continuous-time:}\quad\quad\dot{\delta x}(t)$ $=\dfrac{\partial F}{\partial x}\left(x(t),t\right)\delta x(t).$
@@
All of our experiments are the forward Euler discretizations of continuous-time systems. So, in this section, we work out what happens to the LLE in such a setting. Let’s consider running our continuous-time setting for a length of time $T$.
@@
Note that naively plugging in this discrete-time estimator would therefore result in a different estimate of the LLE depending on the size of the time-step $\Delta t$. However, in most of this paper we still report the naive discrete-time LLE $\lambda_{d}(\Delta t)$ as this is the quantity relates to $\mu$ via $\mathbf{J}$. The exception to our convention is in Table˜1, where we report our estimates $\lambda_{c}$ to better coincide with the intuitions and expectations of readers with backgrounds in continuous time systems. For all systems in Table˜1, we use a step size of $0.01$, and so one can translate from the reported continuous time LLEs to discrete time LLEs by dividing by 100.
Ultimately, dividing by $\Delta t$, which is positive, does not change the sign of $\lambda$, i.e. whether or not the system is predictable or unpredictable.